OpenStudy (anonymous):

The perimeter of a rectangle is 90 ft. Find the dimensions of the rectangle if the ratio of the length to the width is 8: 7. A: 30 ft by 15 ft B: 26 ft by 19 ft C: 24 ft by 21 ft D: 20 ft by 25 ft Express in its simplest form: 6y : 9y A: 2: 3 B: 3: 2 C: 9: 6 D: 6: 9 Please help

hartnn (hartnn):

let the length be 'x' and the width be 'y' so, x/y = 8/7 since perimeter is given, can you find one more equation in x and y ?

OpenStudy (anonymous):

sorry i dont think i understand. when i try work it out i get c but i dont think im right?

hartnn (hartnn):

can you show what have you worked ?

OpenStudy (anonymous):

@LeHa A ratio of 8:7 can be expressed algebraically as 8x:7x. Perimeter = 2l + 2w So your equation would be: 2(8x) + 2(7x) = 90 I think you can solve for x and answer the question.

OpenStudy (anonymous):

@LeHa Do you understand why a ratio of 8:7 can be expressed algebraically as 8x:7x?

OpenStudy (anonymous):

yes now I do thanks

OpenStudy (anonymous):

welcome.